please help me
One family made a trip by car.
They traveled 1312 km in 3 days.
On the 2nd day, it is the number of km from the 1st day.
On the 3rd day, it is 98 km less than the 2nd day.
Find the distance Parcourue every day.

Answer by checkley71(8403)

day 1=x
day 2=x
day 3=x-98
x+x+x-98=1312
3x=1312+98
3x=1410
x=1410/3
x=470 number of km for days 1 & 2.
470-98=372 for day 3
proof
470+470+372=1312 total number of km driven.